OpenStudy (sesca89):

A savings account starts out with $0.50 and doubles every month. Write an equation that models this situation and determine how much money will be in the account in 2 years

OpenStudy (anonymous):

its easy

OpenStudy (anonymous):

you have to multiply .50x2x24

OpenStudy (anonymous):

Because it doubles

OpenStudy (anonymous):

but like .50x2

OpenStudy (anonymous):

then x24

OpenStudy (anonymous):

then 1.00x2

OpenStudy (sesca89):

which is one...

OpenStudy (sesca89):

2....

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

and you have to do that procees 24 times

OpenStudy (anonymous):

i messed up dont do x24

OpenStudy (sesca89):

Well, that's wrong, because it is an exponential growth problem and i need an equation...

OpenStudy (sesca89):

I mean it's right, but it's not what I need

OpenStudy (anonymous):

just do the procces 24 times to dertermine the snaswer

OpenStudy (sesca89):

Yes, but I need an equation too.

OpenStudy (anonymous):

ok

OpenStudy (carlosjose):

hey

OpenStudy (sesca89):

Helpppp

OpenStudy (perl):

y = 0.5*2^x , where x = number of months

OpenStudy (anonymous):

I would go with Y=X exponent 2 x2

OpenStudy (carlosjose):

imma go to english

OpenStudy (sesca89):

x is supposed to be number of years.

OpenStudy (xapproachesinfinity):

it is f(x)=(2^x)/2 like perl said

OpenStudy (perl):

y = 0.5 * 2^(12*x)

OpenStudy (perl):

because it doubles 12 times in a year

OpenStudy (sesca89):

Yes, but wouldn't it be .5(25)^t?

OpenStudy (xapproachesinfinity):

no the amount double each month with the initial value f(0)=0.5

OpenStudy (sesca89):

Listen, this is an exponential growth problem. SO, the equation for it would be A(1 +r)^t A= 0.5, r = 24, and t = # of years (2)

OpenStudy (sesca89):

therefore, it should be 0.5(1+24)^2

OpenStudy (xapproachesinfinity):

this one y=0.5*2^(12x) is an exponential growth as well

OpenStudy (sesca89):

this is confusing....

OpenStudy (xapproachesinfinity):

where did you get r=24? r repesents the rate my dear

OpenStudy (xapproachesinfinity):

in fact you can't say A(1+r)^t this interest formula you know you need to find the interest rate r and the number of compounding

OpenStudy (xapproachesinfinity):

usually as we compound continuously we get the number e^r not 25 like you did

OpenStudy (sesca89):

it's asking for the amount after 2 years. The amount DOUBLES each MONTH 2 x 12(month) = 24

OpenStudy (perl):

r is not 24

OpenStudy (xapproachesinfinity):

why are you making your life harder, while the problem didn't ask you anything about interest ?

OpenStudy (perl):

If something doubles that means r = 100% , or in decimal r = 100/100 = 1

OpenStudy (xapproachesinfinity):

yes @perl has given you the formula for getting that \(y=0.5*(2)^{12x}\)

OpenStudy (sesca89):

Listen, its not asking about intrest, it's asking about what the amount will be in 2 years aka 24 months

OpenStudy (xapproachesinfinity):

I know! you are the one who is thinking about a different approach to the problem

OpenStudy (perl):

y = A ( 1 + r)^t here A = 0.5 r = 1 t = variable in months y = 0.5 ( 1 + 1) ^t and then you said t is in years , so i changed it y = 0.5 ( 1 + 1) ^(12t)

OpenStudy (sesca89):

Because this is the approach I am being taught

OpenStudy (perl):

r is the rate it increases per time period. it increases by 100% or by a factor of 1

OpenStudy (xapproachesinfinity):

hmm, I agree with perl

OpenStudy (sesca89):

you guys are probably right, I'm just asking if the equation, .5(25)^t is right

OpenStudy (perl):

no, it isn't

OpenStudy (perl):

for the reasons given above :)

OpenStudy (sesca89):

...

OpenStudy (sesca89):

Then what would r be?

OpenStudy (perl):

r equals 1

OpenStudy (sesca89):

Wait, just tell me the total answer for the problem.

OpenStudy (sesca89):

Is the total $312.50?